An American ski jumper, has a mass of 59.6 kg. He is moving at a speed of 23.4 m/s at a height of 44.6 meters above the ground. Determine the total mechanical energy of the jumper.


a. 4.42 x 104 J                  

b. 4.24 x 10 -4 J                 

c. 4.24 x 10 3 J                    

d. 42. 24 x 104 J

 


1
Expert's answer
2022-02-25T11:17:27-0500

"A=m(gh+\frac{v^2}2)=4.24\cdot 10^4~J."